Katy invests a total of $24,000 in two accounts paying 7% and 13% simple interest, respectively. How much was invested in each account if, after one year, the total interest was $2,790.00.
A) Enter an equation that uses the information as it is given that can be used to solve this problem. Use
x
as your variable to represent the amount of money invested in the account paying 7% simple interest.
Equation:

B) The answers are:
$
was invested at 7% and
$
was invested at 13%.

Found 3 solutions by stanbon, mananth, ikleyn:
Answer by stanbon(75887)

Katy invests a total of $24,000 in two accounts paying 7% and 13% simple interest, respectively. How much was invested in each account if, after one year, the total interest was $2,790.00.
-----
Int + Int = Int
0.07x + 0.13(24000-x) = 2790
------
7x + 13*24000 - 13x = 279000
-6x = -33000
x = $5500 (amt. invested at 7%
24000-x = $18500 (amt. invested at 13%
-----------
Cheers,
Stan H.
--------------

Answer by mananth(16946)

Katy invests a total of $24,000
account paying 7% ....x
amount paying 13% ,....24000-x
7%x+13%(24000-x) = 2790
multiply equation by 100
7x+13(24000-x) = 279000
7x+312000-13x=279000
simplify
6x= 33000
x=5500 ... at 7%
24000-5500= 18500 at 13%
CHECK
18500*13% +5500*7% =2790



Answer by ikleyn(52778)

.
If you want to learn this subject, look into the lesson
    - Using systems of equations to solve problems on investment
in this site.


Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lesson is the part of this online textbook under the topic "Systems of two linear equations in two unknowns".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.